HanAkoSolo (jamierox4ev3r):

Check my work someone, not sure if I'm going about this in the correct manner

HanAkoSolo (jamierox4ev3r):

Give me a second to type out the problem, thank you

HanAkoSolo (jamierox4ev3r):

Simplify the rational expression: \(\Large\frac{10^{2}+29x+21}{x^{2}-25}\times\frac{x+5}{2x+3}\)

HanAkoSolo (jamierox4ev3r):

So far, I've factored the top portion. So \(10x^{2}+29x+21\) should turn into (5x+7)(2x+3) while \(x^{2}-25\) should be (x-5)(x+5)

HanAkoSolo (jamierox4ev3r):

*factored both the numerator and denominator of the first part of the expression

HanAkoSolo (jamierox4ev3r):

So I tried cross-multiplying, but that brought me more complications...however, I noticed a few things \(\huge\frac{(5x-7)(2x+3)}{(x-5)(x+5)}\times\frac{x+5}{2x+3}\) \(\huge\frac{(5x-7)\color{red}{(2x+3)}}{(x-5)\color{red}{(x+5)}}\times\frac{\color{red}{x+5}}{\color{red}{2x+3}}\) Can the things in red cancel out? And would that lead me with a final answer?

HanAkoSolo (jamierox4ev3r):

^This is where there are some clear holes in my recollection of simplifying. Someone clarify my thinking, thanks

jimthompson5910 (jim_thompson5910):

you are correct. The `2x+3` terms will divide to 1 and effectively go away since 1*(any number) = same number the same applies to the `x+5` terms

HanAkoSolo (jamierox4ev3r):

so \(\huge\frac{5x+7}{x-5}\) would be final answer? I mean, I'm not sure if that could be simplified any further, but I don't think so.

jimthompson5910 (jim_thompson5910):

that's as simplified as it gets

HanAkoSolo (jamierox4ev3r):

Hmm alright. Thank you for your time :)

jimthompson5910 (jim_thompson5910):

you're welcome